May I kindly ask about calculation WACC and explain a little bit that I would like to understand how. It has current capital and new capital determine.

The comic book publishing group (CBPG) has a capital structure of $ 12 million in bonds, paying a 5% coupon, $ 5 million in preferred stock, a face value of $ 35 per share and an annual dividend of $ 1.75 per share. The company's common stock has a book value of $ 6 million. 10% common stock and the company's marginal tax rate is 33%.

The company wants to add another $ 10 million by public debt offering (corporate bond) but bonds at a coupon rate of 4% due to favorable interest rates.
Answer #1
Old capital structure Weightage Formula Weightage Post tax cost Weightage x Post tax cost
Bonds 12 =12/23 52.17% 3.35% 1.75%
Preferred stock 5 =5/23 21.74% 5.00% 1.09%
Common stock 6 =6/23 26.09% 10.00% 2.61%
Total 23 100.00% WACC 5.44%
Post tax cost of bond = pre-tax cost of bond * (1- tax rate)
Post tax cost of bond = 5%*(1-33%)
Post tax cost of bond = 3.35%
Preferred stock cost = Preferred dividend / Preferred stock price
Preferred stock cost = 1.75/35
Preferred stock cost = 5.00%
New capital structure Weightage Formula Weightage Post tax cost Weightage x Post tax cost
Bonds 12 =12/33 36.36% 3.35% 1.22%
Preferred stock 5 =5/33 15.15% 5.00% 0.76%
Common stock 6 =6/33 18.18% 10.00% 1.82%
Public debt 10 =10/33 30.30% 2.68% 0.81%
Total 33 100.00% WACC 4.61%
Post tax cost of bond = pre-tax cost of bond * (1- tax rate)
Post tax cost of bond = 5%*(1-33%)
Post tax cost of bond = 3.35%
Preferred stock cost = Preferred dividend / Preferred stock price
Preferred stock cost = 1.75/35
Preferred stock cost = 5.00%
Post tax cost of public debt = pre-tax cost of debt * (1- tax rate)
Post tax cost of public debt = 4% * (1- 33%)
Post tax cost of public debt = 2.68%
Current capital structure WACC 5.44%
New capital structure WACC 4.61%